find the first five values of the recursive function
f(1)= 5; f(n) = f(n-1) + n

Answer by stanbon(57219)

find the first five values of the recursive function
-------------------------------
f(1)= 5;
f(n) = f(n-1) + n
f(1) = 5
f(2) = 5 + 2 = 7
f(3) = 7 + 3 = 10
f(4) = 10 + 4 = 14
f(5) = 14 + 5 = 19
-----------------------
Cheers,
Stan H.
===========